OpenStudy (anonymous):

-5 × 9 ------- -21 + 6

OpenStudy (pfenn1):

\[\frac{-5 × 9}{-21 + 6}=\frac{-45}{-15}=?\]

OpenStudy (anonymous):

-5/-3?

OpenStudy (pfenn1):

Check your math again, your fraction isn't correct, and the negatives cancel out.

OpenStudy (anonymous):

ohhhh, ok so its 5/3...

OpenStudy (pfenn1):

No, that is not the answer. How did you get 5/3?

OpenStudy (anonymous):

oh, wait. i know what i did wrong, i meant 3

OpenStudy (pfenn1):

Correct!